Given a normal population whose mean is 50 and whose standard deviation is 5, find the probability that a random sample of a. 4 has a mean between 49 and 52. b. 16 has a mean between 49 and 52. c. 25 has a mean between 49 and 52


1
Expert's answer
2021-06-25T08:59:19-0400

"\mu=50 \\

\sigma=5"

a. n=4

"P(49<\bar{x}<52) = P(\bar{x}<52) -P(\bar{x}<49) \\

=P(Z< \frac{52-50}{5 / \sqrt{4}}) -P(Z< \frac{49-50}{5 / \sqrt{4}}) \\

=P(Z< 0.8) -P(Z< -0.4) \\

= 0.7881 - 0.3445 \\

=0.4436"

b. n=16

"P(49<\bar{x}<52) = P(\bar{x}<52) -P(\bar{x}<49) \\

=P(Z< \frac{52-50}{5 / \sqrt{16}}) -P(Z< \frac{49-50}{5 / \sqrt{16}}) \\

=P(Z< 1.6) -P(Z< -0.8) \\

= 0.9452 - 0.2118 \\

=0.7334"

c. n=25

"P(49<\bar{x}<52) = P(\bar{x}<52) -P(\bar{x}<49) \\

=P(Z< \frac{52-50}{5 / \sqrt{25}}) -P(Z< \frac{49-50}{5 / \sqrt{25}}) \\

=P(Z< 2) -P(Z< -1) \\

= 0.9772 - 0.1586 \\

=0.8186"